### v1

# Explicit Multimodal Graph Modeling for Human-Object Interaction Detection

Abstract
Transformer-based methods have recently become the prevailing approach for Human-Object Interaction (HOI) detection. However, the Transformer architecture does not explicitly model the relational structures inherent in HOI detection, which impedes the recognition of interactions. In contrast, Graph Neural Networks (GNNs) are inherently better suited for this task, as they explicitly model the relationships between human-object pairs. Therefore, in this paper, we propose Multimodal Graph Network Modeling (MGNM) that leverages GNN-based relational structures to enhance HOI detection. Specifically, we design a multimodal graph network framework that explicitly models the HOI task in a four-stage graph structure. Furthermore, we introduce a multi-level feature interaction mechanism within our graph network. This mechanism leverages multi-level vision and language features to enhance information propagation across human-object pairs. Consequently, our proposed MGNM achieves state-of-the-art performance on two widely used benchmarks: HICO-DET and V-COCO. Moreover, when integrated with a more advanced object detector, our method demonstrates a significant performance gain and maintains an effective balance between rare and non-rare classes.
Introduction
Figure 1: Comparison of existing Transformer-based methods and our GNN-based method. Transformer-based methods typically perform feature extraction prior to matching without explicit relational modeling, making it difficult for them to identify complex interactions. In contrast, our GNN-based method firstly constructs human-object pairs via a general matching mechanism and then applies multi-level feature interaction mechanism to enable explicit GNN-based relational reasoning. H, O, and f denote human, object, and multimodal features, respectively.
Human-object interaction detection is a task that aims to localize human-object pairs and simultaneously infer the interactions between them. This task represents a more profound level of visual scene understanding than standard object detection. A detected HOI instance is represented as a triplet in the form of $\langle\text{human, action, object}\rangle$. Recently, HOI detection has garnered significant attention from the computer vision community, owing to its extensive applications in downstream tasks such as action recognition (Kong and Fu 2022), image caption (Hossain et al. 2018), and visual scene understanding (Ma et al. 2024a).
Similar to object detectors, HOI detectors can be categorized into two primary types: one-stage and two-stage methods. Recent one-stage methods usually employ the Transformer architecture. Specifically, these methods often adapt DETR-like frameworks (Carion et al. 2020) to detect human-object pairs and predict the interactions in parallel via Transformer decoders (Liao et al. 2022; Zhou et al. 2022). However, compared with two-stage methods, one-stage methods cannot easily leverage the features from object detectors, potentially limiting their performance.
Two-stage methods typically follow a distinct training paradigm, which generally involves fine-tuning a pre-trained object detector and subsequently freezing its parameters (Zhang, Campbell, and Gould 2022; Zhang et al. 2023). This paradigm allows the interaction predictor to focus solely on interaction prediction, thereby improving training efficiency. As illustrated in Figure 1(a), recent two-stage methods usually adopt Transformer to extract the visual features. However, the attention mechanism inherent in the Transformer updates features via a weighted-sum aggregation, which does not explicitly model the relational structure in HOI detection and thus hinders further performance gains. Motivated by this limitation, we propose a multimodal graph network framework designed to explicitly model the HOI detection task, which more closely aligns with the human perception of HOI.
As illustrated in Figure 1(b), our GNN-based method begins by matching detected human and object instances to construct explicit relational connections for each human-object pair. Notably, as observed in prior studies (Zhang, Campbell, and Gould 2021; Park, Park, and Lee 2023), relying solely on the visual features of human and object nodes often fails to capture fine-grained contextual cues. Inspired by the success of Vision Language Models (VLMs) (Lei, Wang, and Tan 2024; Guo et al. 2024), we leverage CLIP (Radford et al. 2021) to generate rich multimodal features, thereby enhancing feature interaction between each human-object pair. Benefiting from the sufficient prior knowledge of CLIP, our method not only achieves a significant performance gain but also effectively mitigates the performance disparity between rare and non-rare classes, which is a persistent challenge for previous methods (Zhang, Campbell, and Gould 2021, 2022; Liao et al. 2020). Moreover, we introduce a Multi-level Feature Interaction (MFI) mechanism designed to further utilize these multimodal features and enhance the exchange of contextual information. This mechanism leverages both low-level and high-level vision and language features to intensify the interaction between human-object pairs, thereby enriching the graph network with more fine-grained semantic information.
In light of the preceding analysis, we propose Multimodal Graph Network Modeling (MGNM), a novel and efficient two-stage framework for HOI detection. Our MGNM achieves state-of-the-art performance on two widely used benchmarks: HICO-DET (Chao et al. 2018) and V-COCO (Gupta and Malik 2015). Furthermore, when integrated with a stronger object detector, our method demonstrates a significant performance gain and maintains an effective balance between rare and non-rare classes. The effectiveness of each proposed component is validated through extensive ablation studies. The contributions of this work are as follows:
  • We propose a novel multimodal graph network framework that explicitly models the relational structure in HOI detection, addressing the limitations of Transformer-based HOI detectors that rely on implicit attention mechanisms.
  • We design a multi-level feature interaction mechanism that effectively fuses low-level and high-level semantic cues across modalities, substantially enhancing the exchange of fine-grained contextual information within the graph structure.
  • Extensive experiments on HICO-DET and V-COCO benchmarks demonstrate that our method not only achieves state-of-the-art performance but also maintains an effective balance between rare and non-rare classes.
Related Work
One-stage Methods
One-stage methods are characterized by their parallel detection of human and object instances and prediction of their corresponding interactions. Early approaches typically relied on heuristics such as human keypoints (Liao et al. 2020) or predefined union boxes (Kim et al. 2020). More recently, Transformer-based architectures have become the dominant paradigm for one-stage HOI detection. For instance, Chen et al. (Chen et al. 2021) employ two separate Transformer decoders to predict human and object instances and their interactions in parallel. Similarly, Yang et al. (Yang et al. 2023) leveraged VLMs to extract fine-grained semantic features, which were then processed by two sub-decoders to simultaneously predict HOI triplets. Nevertheless, one-stage methods are often encumbered by significant memory usage and slow convergence speed.
Two-stage Methods
Two-stage methods first employ an off-the-shelf object detector to identify human and object instances, and subsequently predict interactions between the matched human-object pairs. Recent two-stage methods often use Transformer encoders to extract features or enhance attention mechanisms (Zhang, Campbell, and Gould 2022; Zhang et al. 2023). Another line of work (Guo et al. 2024; Lei et al. 2023; Lei, Wang, and Robby T. 2024) has incorporated VLMs for feature refinement, leading to notable accuracy improvements. However, while the Transformer excels as a powerful feature extractor, it does not offer an explicit relational modeling framework for the HOI detection task (Antoun and Asmar 2023). To address this gap, our method employs a multimodal graph network framework instead. Our work is situated within a growing line of GNN-based HOI detectors. For instance, SCG (Zhang, Campbell, and Gould 2021) utilized a GNN to facilitate information exchange between human and object nodes, while ViPLO (Park, Park, and Lee 2023) incorporated features from the CLIP image encoder and ViTPose (Xu et al. 2022) to enrich this interaction. Nevertheless, a key limitation of these methods is their primary reliance on visual features, which overlooks the rich semantic cues offered by multimodal features, thereby limiting their ultimate performance.
Method
Figure 2: Overview of our MGNM framework.The proposed MGNM framework initiates by employing the DETR model to detect human and object instances, which are subsequently matched to form candidate human-object pairs. The features of these pairs are then processed through a four-stage multimodal graph network, consisting of the spatial, visual, textual, and interaction stages. (1) Spatial Stage: For each candidate pair, low-level spatial features derived from their bounding boxes are used to initialize pairwise representations. Simultaneously, these spatial features are utilized to construct a weighted adjacency matrix that regulates feature interactions in the subsequent stages. (2) Visual Stage: High-level visual semantic features are extracted using the CLIP image encoder. These features further enrich the interactions among human-object pairs by incorporating detailed visual context. (3) Textual Stage: The CLIP text encoder is adopted to obtain semantic embeddings for the corresponding human and object categories, providing complementary textual cues for each pair. (4) Interaction Stage: In the final stage, the model captures high-level interaction features between human-object pairs, facilitating more effective relational reasoning within the graph structure. Finally, the refined pairwise representations are utilized as queries in a Transformer decoder, which attends to global image features from the backbone (serving as keys and values), to predict the final HOI triplets.
This section provides a detailed description of our method. We first elaborate on our multi-level feature interaction mechanism, which is structured into four distinct stages. Subsequently, we present the overall framework of MGNM. An overview of our method is illustrated in Figure 2.
Following the standard two-stage paradigm for HOI detection, we assume that a given image contains $m$ detected humans and $n$ detected objects, resulting in $m\times n$ candidate human-object pairs. Here, we denote $\mathbf{N}\in\mathbb{R}^{(m+n)\times d}$ as the human and object instances, which serves as the node representations in the GNN. The corresponding human and object sets within the human-object pairs are represented as $\mathbf{H}\in\mathbb{R}^{(m\times n)\times d}$ and $\mathbf{O}\in\mathbb{R}^{(m\times n)\times d}$, respectively. $d$ is the dimension of the nodes. Additionally, the features of human-object pairs are denoted as $\mathbf{E}\in\mathbb{R}^{(m\times n)\times 2d}$, which can be interpreted as the edge embeddings within the GNN framework.
Multi-level Feature Interaction Mechanism
In this paper, we propose a novel Multi-level Feature Interaction (MFI) mechanism. This mechanism utilizes the low-level and high-level vision and language features to enhance the interactive modeling of human-object pairs. As illustrated in Figure 2, the MFI mechanism is structured into four sequential stages: Spatial, Visual, Textual, and Interaction. The details of each stage are elaborated below.
Spatial Stage
The spatial stage is designed to exploit geometric relationships to enhance interactions between human-object pairs. Specifically, the spatial features comprise bounding box coordinates, box areas, Intersection over Union (IoU), and other related metrics (Zhang, Campbell, and Gould 2021). These features encode the fundamental geometric configurations of each human-object pair and serve as a source of low-level visual information. Moreover, since these are handcrafted features, their computation incurs minimal overhead for the model (Zhang, Campbell, and Gould 2022). Upon extracting the spatial features, we utilize a Multimodality Fusion (MMF) mechanism (Zhang et al. 2023) to integrate them with the initial human and object features, thereby generating the initial representation for each human-object pair.
$\mathbf{E}$ $=\text{MMF}(\mathbf{H}\oplus\mathbf{O},\mathbf{S}),$ (1)
where $\oplus$ means the concatenation operation and $\mathbf{S}$ means the spatial features. In addition, based on the spatial features and corresponding human and object features, the weight matrix $\mathbf{W}_{G}$ for the multimodal graph network is attained with the multi-branch fusion (MBF) mechanism (Zhang, Campbell, and Gould 2021).
$\mathbf{W}_{G}$ $=\text{Linear}(\text{MBF}(\mathbf{H}\oplus\mathbf{O},\mathbf{S})),$ (2)
where the linear layer is used to align the dimensions.
Visual Stage
The visual stage leverages the CLIP to incorporate high-level vision features. Specifically, we employ the CLIP image encoder to extract high-level visual semantic features from the given image. These visual features are then used to pass the high-level visual semantic information to the humans and objects as follows:
$\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}(\mathbf{V},\mathbf{N})$ $=\sigma(\mathbf{W}_{G}\odot\text{MBF}_{v}(\mathbf{V},\mathbf{N})),$ (3)
where $\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}$ denotes the message passing function responsible for propagating visual information to the human and object node features. $\mathbf{V}$ represents the visual features from CLIP. $\sigma$ is ReLU activation function and $\odot$ is element-wise multiplication. $\text{MBF}_{v}$ is a subclass of the MBF mechanism, tasked with integrating visual and node features. Finally, the human and object node features are updated using a layer normalization step, as formulated below:
$\mathbf{N}$ $=\text{LN}(\mathbf{N}+\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}(\mathbf{V},\mathbf{N})),$ (4)
where LN means the layer normalization.
Textual Stage
The textual stage is dedicated to generating textual embeddings for each node, providing a source of low-level language features. As illustrated in Figure 2, we adopt the prompt template “a photo of a $\langle\text{object}\rangle$”, where the $\langle\text{object}\rangle$ comes from the detection results. In this paper, for brevity, we denote the humans and objects as the whole node set $\mathbf{N}$, and the corresponding textual features are assumed to be matched. The process is formulated as follows:
$\text{M}_{\mathcal{T}\rightarrow\mathcal{N}}(\mathbf{T},\mathbf{N})$ $=\sigma(\mathbf{W}_{G}\odot\text{MBF}_{t}(\mathbf{T},\mathbf{N})),$ (5)
where $\text{M}_{\mathcal{T}\rightarrow\mathcal{N}}$ means the message passing function responsible for propagating textual information to the human and object node features. $\mathbf{T}$ means the textual features from CLIP. $\text{MBF}_{t}$ is a subclass of the MBF mechanism. Similarly, we also use layer normalization step when updating the human and object features.
$\mathbf{N}$ $=\text{LN}(\mathbf{N}+\text{M}_{\mathcal{T}\rightarrow\mathcal{N}}(\mathbf{T},\mathbf{N})).$ (6)
Interaction Stage
To incorporate high-level interaction semantic cues, we again employ the CLIP text encoder. This time, it is used to generate interaction-centric semantic features, which constitute high-level language features. The prompt template “a photo of a person interacting with $\langle\text{object}\rangle$” is adopted, which aims to boost the feature interaction between human-object pairs. This process is formulated as follows:
$\mathbf{E}$ $=\text{LN}(\mathbf{E}+\mathbf{I}),$ (7)
where $\mathbf{I}$ denotes the interaction features of the human-object pairs. As a preliminary step, the human-object pair features are first enriched by integrating the features of their corresponding human and object nodes as follows:
$\mathbf{E}$ $=\text{LN}(\mathbf{E}+\mathbf{H}\oplus\mathbf{O}).$ (8)
Multimodal Graph Network Modeling
Building upon the MFI mechanism detailed above, this section presents the complete architecture of the proposed MGNM framework.
Algorithm 1 Computation Process of Interaction Prediction
Input: $\mathbf{N},\mathbf{H},\mathbf{O}$, Image
Output: $\text{Decoder}(\mathbf{Q},\mathbf{K},\mathbf{V})$
[…]
2:for $i=1:T$ do
3:  $\mathbf{W}_{G}=\text{Linear}(\text{MBF}(\mathbf{H}\oplus\mathbf{O},\mathbf{S}))$
4:  $\mathbf{N}=\text{LN}(\mathbf{N}+\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}(\mathbf{V},\mathbf{N}))$
5:  $\mathbf{N}=\text{LN}(\mathbf{N}+\text{M}_{\mathcal{T}\rightarrow\mathcal{N}}(\mathbf{T},\mathbf{N}))$
6:  $\mathbf{E}=\text{LN}(\text{LN}(\mathbf{E}+\mathbf{H}\oplus\mathbf{O})+\mathbf{I})$
7:end for
[…]
9:return $\text{Decoder}(\mathbf{Q},\mathbf{K},\mathbf{V})$
As illustrated in Figure 2, the process begins with an off-the-shelf object detector, which extracts bounding boxes and appearance features. Based on these detection results and a matching strategy, we construct human and object instances $\mathbf{N}$, along with human features $\mathbf{H}$ and object features $\mathbf{O}$. At the core of our framework lies the MFI mechanism, as detailed in Algorithm 1 (Lines 1–7). The process commences by constructing the initial human-object pair features, which are then iteratively refined through the proposed four-stage pipeline. Following (Park, Park, and Lee 2023), we set the number of iterative steps $T$ to 2. Notably, in the visual and textual stages, we leverage visual and textual features to update the human and object instances, thereby infusing multimodal information into the node encodings. In addition, since the message passing is guided by a weight matrix computed from the relational structure among human-object pairs, this implicitly promotes feature interaction between pairs. In contrast, the spatial and interaction stages are explicitly designed to enhance feature communication between human-object pairs. Finally, a decoder is employed to incorporate contextual semantic information and produce the final predictions. In this decoder, the human-object pair features act as queries, while the global features extracted from the backbone serve as keys and values.
Experiments
HICO-DET V-COCO
Method Backbone Default Known Object
Full Rare Non-Rare Full Rare Non-Rare $AP^{S1}_{role}$ $AP^{S2}_{role}$
One-stage methods
UnionDet (2020) R50-F 17.58 11.72 19.33 19.76 14.68 21.27 47.5 56.2
PPDM (2024) Swin-B-F 27.59 18.07 30.44 29.11 18.85 32.71 - -
AS-Net (2021) R50 28.87 24.25 30.25 31.74 27.07 33.14 53.9 -
FGAHOI-S (2024b) Swin-T 29.94 22.24 32.24 32.48 24.16 34.97 60.5 61.2
DOQ (2022) R50+C 33.28 29.19 34.50 - - - 63.5 -
GEN-VLTK (2022) R50+C 33.75 29.25 35.10 36.78 32.75 37.99 62.4 64.4
DiffHOI-S (2023) R50+C+SD 34.41 31.07 35.40 37.31 34.56 38.14 61.1 63.5
FGAHOI-L (2024b) Swin-L 37.18 30.71 39.11 38.93 31.93 41.02 - -
DiffHOI-L (2023) Swin-L+C+SD 40.63 38.10 41.38 43.14 40.24 44.01 65.7 68.2
Two-stage methods
SCG (2021) R50-F 29.26 24.61 30.65 32.87 27.89 34.35 54.2 60.9
PPDM++ (2024) Swin-B-F 30.10 23.73 32.00 31.80 24.93 33.85 - -
UPT (2022) R101 32.31 28.55 33.44 35.65 31.60 36.86 60.7 66.2
ADA-CM (2023) R50+C 33.80 31.72 34.42 - - - 60.7 66.2
ILCN (2025) Swin-T 33.80 29.83 34.99 - - - 63.4 65.3
PViC w/ DETR (2023) R50 34.69 32.14 35.45 38.14 35.38 38.97 62.8 67.8
HOIGen (2024) R50+C 34.84 34.52 34.94 - - - - -
ViPLO (2023) R101-D+C+VP 34.95 33.83 35.28 38.15 36.77 38.56 60.9 66.6
TKCE (2025) R50+C 35.11 34.35 35.34 - - - - -
LAIN (2025) R50+C 36.02 35.70 36.11 - - - 63.4 65.3
MP-HOI-S (2024) R50+C+SD 36.50 35.48 36.80 - - - 66.2 67.6
PViC w/ $\mathcal{H}$-DETR (2023) Swin-L 44.32 44.61 44.24 47.81 48.38 47.64 64.1 70.2
MP-HOI-L (2024) Swin-L+C+SD 44.53 44.48 44.55 - - - - -
MGNM R50+C 36.68 34.10 37.45 39.99 37.15 40.84 62.1 67.5
MGNM w/ $\mathcal{H}$-DETR Swin-L+C 46.54 48.09 46.07 49.78 50.82 49.47 64.3 70.5
Table 1: Performance comparison in terms of mAP (%) on the HICO-DET and V-COCO dataset. To ensure a fair comparison, we report both the performance and the corresponding feature backbone for each method. In each evaluation metric, the best result is marked in bold and the second-best result is underlined. In this paper, R50, R101, R50-F and R101-D denotes ResNet50, ResNet101, ResNe50-FPN and ResNet101-DC5 model, respectively. A similar notation applies to Swin Transformer (Liu et al. 2021). C, SD, and VP mean CLIP, Stable Diffusion, and ViTPose models, respectively. We note that MP-HOI (Yang et al. 2024) utilizes additional training data (Magic-HOI and SynHOI), which should be considered when comparing results.
This section presents a series of experiments designed to evaluate the effectiveness of our proposed method. First, we briefly introduce the experimental settings. Next, we present a comprehensive comparison of our method against existing state-of-the-art methods. Subsequently, the ablation study of our method is conducted. Finally, we rethink and analyze the Rare-Non-rare bias in the HOI detection field and provide some promising suggestions.
Experimental Settings
Datasets
Our experiments are conducted on two widely used public benchmarks: HICO-DET and V-COCO. HICO-DET contains 38,118 training images and 9,658 test images. It includes 80 object classes from the MS COCO dataset (Lin et al. 2014) and 117 action classes, which combine to form 600 types of HOI instances. V-COCO, a smaller-scale subset of the MS COCO dataset, contains 2,533 training, 2,867 validation, and 4,946 test images, encompassing 80 object classes and 29 action classes. Given its larger scale and closer alignment with real-world data distributions, our primary experimental focus is on the HICO-DET dataset.
Evaluation Metrics
Following the standard evaluation protocols in (Park, Park, and Lee 2023), we report the mean Average Precision (mAP) for evaluation. An HOI triplet is considered a true positive when the IoU of the detected human and object is larger than 0.5 and the action is predicted accurately. For the HICO-DET dataset, we report performance under two standard evaluation settings: Default and Known Object. Each setting has three different subsets: (1) all 600 HOI classes (Full), (2) 138 HOI classes with less than 10 training instances (Rare), and (3) 462 HOI classes with 10 or more training instances (Non-Rare). For the V-COCO dataset, we evaluate performance based on two scenarios: Scenario 1, which requires correct prediction of occluded object bounding boxes, and Scenario 2, which does not.
Implementation Details
For HICO-DET and V-COCO datasets, we use the fine-tuned DETR and $\mathcal{H}$-DETR (Jia et al. 2023) weights provided by Zhang et al. (Zhang et al. 2023). During training, following the commonly used training paradigm of two-stage methods, the weights of the object detector are frozen, while only the interaction predictor is trained. Meanwhile, the focal loss (Lin et al. 2017) is employed as the loss function. Following (Lei et al. 2023), we also incorporate the adapter into the CLIP module. The model is optimized using the AdamW optimizer with an initial learning rate of $10^{-4}$. All models are trained for 30 epochs on 4 NVIDIA A100 GPUs, using a total batch size of 32. Unless otherwise specified, all other hyperparameters adhere to the settings established in (Zhang et al. 2023; Lei et al. 2023).
Comparison with State-of-the-art Methods
Table 1 provides a comprehensive performance comparison between our proposed MGNM and current state-of-the-art methods. For clarity, the competing methods are grouped into one-stage and two-stage categories. On the HICO-DET dataset, MGNM outperforms all other one-stage and two-stage approaches that utilize a ResNet backbone. Specifically, in the Default Full setting, our method achieves a 2.27 mAP improvement over DiffHOI-S, despite the latter leveraging two VLMs, namely CLIP and Stable Diffusion (Rombach et al. 2021). MGNM also surpasses the leading two-stage approach, MP-HOI-S, by 0.18 mAP in the Default Full setting, even though MP-HOI-S incorporates both additional VLMs and extra training data. Moreover, MGNM exceeds the performance of the recent GNN-based ViPLO by 1.73 mAP in the Default Full setting, notwithstanding ViPLO’s adoption of the supplementary ViTPose-L model (Xu et al. 2022). In the Known Object setting, MGNM consistently outperforms all ResNet-based counterparts across all three subsets by a substantial margin. Notably, our ResNet50-based MGNM even surpasses FGAHOI-L, which utilizes a more powerful Swin-L backbone, by 1.06 mAP in the Full setting. Collectively, these results demonstrate the effectiveness of our multimodal graph network and its explicit modeling strategy for HOI detection. On the V-COCO dataset, MGNM also achieves highly competitive performance among ResNet-based methods. We note that MP-HOI-S obtains a higher $AP_{role}^{S1}$, which we attribute to its use of additional training data. This observation suggests that GNN-based approaches, including MGNM, could further benefit from large-scale pre-training, highlighting a promising direction for future research.
Methods Full Rare Non-Rare
w/o Spatial Stage 36.00 34.68 36.39
w/o Visual Stage 35.32 32.66 36.12
w/o Textual Stage 36.16 33.81 36.86
w/o Interaction Stage 35.61 33.37 36.28
vanilla 36.68 34.10 37.45
Table 2: Effectiveness of each component within our MFI mechanism. w/o means “without”.
Furthermore, to further demonstrate the scalability of our approach, we compare the performance of MGNM equipped with a more powerful object detector against other state-of-the-art methods that utilize similarly strong backbones. As shown in Table 1, MGNM integrated with $\mathcal{H}$-DETR achieves substantial performance improvements over the competing methods. For example, on the HICO-DET Default Full setting, our method outperforms PViC with $\mathcal{H}$-DETR by 2.22 mAP. Notably, although both DiffHOI-L and MP-HOI-L leverage two VLMs each, our single-VLM approach achieves significant gains of 5.91 and 2.01 mAP over these methods, respectively. These results underscore not only the scalability of our framework but also the effectiveness of our multimodal graph network design.
Ablation Study
(a) carry sports ball
(b) ride snowboard
(c) drive boat
(d) Visualization for c
(e) Visualization for c
Figure 3: Qualitative results on the HICO-DET test set. Figures 3(a) and 3(b) illustrate successful predictions across a variety of challenging scenarios. In contrast, Figure 3(c) presents a representative failure case, wherein the model generates a false positive triplet. Specifically, it incorrectly infers that both individuals are driving the boat, whereas only the person on the right is performing this action. To further investigate this failure case, Figures 3(d) and 3(e) visualize the corresponding attention maps, highlighting the model’s regions of focus that led to this prediction.
Methods Backbone mAP $\uparrow$ Time $\downarrow$ Memory $\downarrow$
SCG R101 29.26 182 ms 7,702 MiB
ViPLO R101-D+C+VP 34.95 166 ms 10,224 MiB
MGNM R50+C 36.68 125 ms 2,946 MiB
Table 3: Comparison of mAP, inference time, and GPU memory on the HICO-DET test set. We select two prominent GNN-based baselines for this comparison: SCG and ViPLO. It is important to note that the official implementations of both SCG and ViPLO execute object detection and interaction prediction as two separate, sequential processes. Therefore, to ensure a fair comparison, our reported time and memory for these baselines encompass the complete process. All experiments are conducted with a batch size of 1 on a single NVIDIA A100 GPU. Time refers to the elapsed time required for processing one image.
In this subsection, we conduct ablation studies to analyse the effectiveness of our proposed components and the efficiency of our MGNM framework. All the experiments are conducted on the HICO-DET test set under Deafult setting.
Components Analysis
We first analyze the contribution of each of the four stages within our MFI mechanism, which collectively process low-level and high-level multimodal features. As shown in Table 2, removing the low-level feature stages, namely the Spatial (row 1) and Textual (row 3) stages, results in performance drops of 0.68 and 0.52 mAP, respectively. This demonstrates that these low-level spatial and textual features provide a valuable supplement to the foundational human-object pair representations. Ablating the high-level feature stages, the Visual (row 2) and Interaction (row 4) stages, induces a more substantial performance degradation, with drops of 1.36 and 1.07 mAP, respectively. These decreases are considerably larger than those observed in low-level stages. This finding suggests that the high-level features are more critical to our model’s predictive power, an observation that aligns with the notion that GNNs are inherently biased towards modeling high-level structural information (Antoun and Asmar 2023). Interestingly, we observe a counter-intuitive phenomenon in row 1: ablating the spatial stage improves performance on the rare classes. This suggests that low-level vision features may inadvertently amplify the model’s existing bias towards more frequent, non-rare classes. Nevertheless, considering their substantial positive impact on the Full and Non-Rare settings, these spatial features remain an indispensable component for overall model performance.
Efficiency Analysis
This subsection presents an efficiency analysis of MGNM in comparison to two related GNN-based methods, focusing on inference time and GPU memory usage. As reported in Table 3, both SCG and ViPLO decouple the object detection and interaction prediction processes, whereas our method integrates these stages into a unified end-to-end pipeline. To ensure a fair comparison, we report the total inference time and GPU memory for SCG and ViPLO, excluding file I/O operations. As shown in Table 3, our approach not only achieves substantially higher mAP but also demonstrates superior efficiency, recording the lowest inference time and GPU memory consumption among the compared methods. These results highlight the strong balance between effectiveness and efficiency achieved by MGNM. It is worth noting that the official object detection scripts provided by the baseline methods appear to be unoptimized with respect to memory usage. When replacing their detection modules with DETR weights, resulting in less GPU consumption than the actual GPU memory usage, the measured occupancy drops to 1,904 MiB for SCG and 4,426 MiB for ViPLO. Even under these conditions, when comprehensively considering both mAP and inference time, our method continues to deliver the best overall effectiveness and efficiency.
Figure 4: Comparison of our method and two related GNN-based methods on rare and non-rare classes.
Rethinking Rare-Non-rare Bias
The challenge of few-shot learning is particularly pronounced in HOI detection, where it manifests as a significant performance disparity between rare and non-rare classes. As evidenced in Table 1, this long-tail distribution problem is severe in many methods. For instance, PPDM and FGAHOI-S exhibit a performance gap of over 10 mAP between their rare and non-rare classes. Similarly, in Figure 4, the GNN-based method SCG also has a significant difference between rare and non-rare classes. However, a key observation from Table 1 is that methods incorporating VLMs substantially mitigate this disparity. This suggests that the rich prior knowledge in VLMs can effectively alleviate the Rare-Non-rare bias. Our own results, illustrated in Figure 4, support this hypothesis: both our MGNM and ViPLO, which leverage CLIP, demonstrate a much narrower performance gap between rare and non-rare classes. Interestingly, ViPLO achieves a better balance than our method. We hypothesize that its use of an additional ViTPose-L model provides another source of rich prior knowledge, playing a similar role like VLMs. Moreover, as shown in Table 1 and Figure 4, an unexpected trend emerges for methods equipped with the $\mathcal{H}$-DETR object detector, these models exhibit a remarkable reversal, where performance on rare classes actually exceeds that on non-rare classes. We posit that this is likely attributable to the unique architectural design of $\mathcal{H}$-DETR, and a deeper investigation into this phenomenon presents a compelling direction for future research. In summary, our analysis indicates that leveraging VLMs and advanced object detectors represents a highly effective strategy for mitigating the Rare-Non-rare bias in HOI detection.
Qualitative Results and Analysis
This section presents a qualitative analysis of our method’s predictions, highlighting both its strengths and current limitations. As illustrated in Figures 3(a) and 3(b), our model demonstrates robustness by accurately predicting HOI triplets even when the target object is small or occluded. Conversely, Figure 3(c) presents a representative failure case involving a false positive prediction. Here, the model incorrectly infers that two individuals are driving the boat, a prediction that contradicts common sense in the real world. To diagnose this error, we visualize the corresponding attention maps in Figures 3(d) and 3(e). The visualization reveals that the model correctly identifies promising regions associated with the drive action for both individuals. However, it lacks the commonsense knowledge to understand that only one person can drive a single boat, leading to the erroneous prediction. Furthermore, attention is diffuse and does not pinpoint fine-grained cues, such as hand-on-wheel placement. This limitation highlights a crucial direction for future work: developing methods to inject real-world commonsense reasoning into HOI detectors to resolve such logical ambiguities.
Conclusion
In this paper, we introduce Multimodal Graph Network Modeling (MGNM), a novel and effective two-stage HOI detection framework. The central contribution of MGNM lies in its explicit modeling of the HOI detection task via the construction of a multimodal graph network. Furthermore, by leveraging Vision Language Models (VLMs), we design a multi-level feature interaction mechanism that systematically fuses low-level and high-level vision and language features, thereby enhancing information propagation within the graph structure. Extensive experiments validate that our approach not only achieves state-of-the-art performance but also delivers superior computational efficiency. In addition, our analysis of the pervasive class imbalance issue provides new insights and practical recommendations for mitigating the Rare-Non-rare bias in future research.

### v2

@@
 Furthermore, we introduce a multi-level feature interaction mechanism within our graph network.
-This mechanism leverages multi-level vision and language features to enhance information propagation across human-object pairs.
-Consequently, our proposed MGNM achieves state-of-the-art performance on two widely used benchmarks: HICO-DET and V-COCO.
+This mechanism leverages multi-level visual and language features to enhance information propagation across human-object pairs.
+Consequently, our proposed MGNM achieves state-of-the-art (SOTA) performance on two widely used benchmarks: HICO-DET and V-COCO.
 Moreover, when integrated with a more advanced object detector, our method demonstrates a significant performance gain and maintains an effective balance between rare and non-rare classes.
-Introduction
+1 Introduction
 Figure 1: Comparison of existing Transformer-based methods and our GNN-based method.
@@
 A detected HOI instance is represented as a triplet in the form of $\langle\text{human, action, object}\rangle$.
-Recently, HOI detection has garnered significant attention from the computer vision community, owing to its extensive applications in downstream tasks such as action recognition (Kong and Fu 2022), image caption (Hossain et al.
-2018), and visual scene understanding (Ma et al.
-2024a).
+Recently, HOI detection has garnered significant attention from the computer vision community [wang2024interaction, Jia2024OrchestratingTS, Yang2025NoMS, Zhang2025HOIDR1RL, Xue2025GuidingHI, 10658050], owing to its extensive applications in downstream tasks such as action recognition [kong2022human, do2024skateformer], image caption [Hossain2018ACS, rotstein2024fusecap], and visual scene understanding [ma2024scene].
 Similar to object detectors, HOI detectors can be categorized into two primary types: one-stage and two-stage methods.
-Recent one-stage methods usually employ the Transformer architecture.
-Specifically, these methods often adapt DETR-like frameworks (Carion et al.
-2020) to detect human-object pairs and predict the interactions in parallel via Transformer decoders (Liao et al.
-2022; Zhou et al.
-2022).
+Recent one-stage methods usually employ the Transformer architecture [Vaswani2017AttentionIA].
+Specifically, these methods often adapt DETR [Carion2020EndtoEndOD] to detect human-object pairs and predict the interactions in parallel via Transformer decoders [Zhou2022HumanObjectID, Jia2024ContextHOISC].
 However, compared with two-stage methods, one-stage methods cannot easily leverage the features from object detectors, potentially limiting their performance.
-Two-stage methods typically follow a distinct training paradigm, which generally involves fine-tuning a pre-trained object detector and subsequently freezing its parameters (Zhang, Campbell, and Gould 2022; Zhang et al.
-2023).
+Two-stage methods typically follow a distinct training paradigm, which generally involves fine-tuning a pre-trained object detector and subsequently freezing its parameters [zhang2022upt, zhang2023pvic, park2023viplo].
 This paradigm allows the interaction predictor to focus solely on interaction prediction, thereby improving training efficiency.
@@
 As illustrated in Figure 1(b), our GNN-based method begins by matching detected human and object instances to construct explicit relational connections for each human-object pair.
-Notably, as observed in prior studies (Zhang, Campbell, and Gould 2021; Park, Park, and Lee 2023), relying solely on the visual features of human and object nodes often fails to capture fine-grained contextual cues.
-Inspired by the success of Vision Language Models (VLMs) (Lei, Wang, and Tan 2024; Guo et al.
-2024), we leverage CLIP (Radford et al.
-2021) to generate rich multimodal features, thereby enhancing feature interaction between each human-object pair.
-Benefiting from the sufficient prior knowledge of CLIP, our method not only achieves a significant performance gain but also effectively mitigates the performance disparity between rare and non-rare classes, which is a persistent challenge for previous methods (Zhang, Campbell, and Gould 2021, 2022; Liao et al.
-2020).
+Notably, as observed in prior studies [zhang2021scg, park2023viplo], relying solely on the visual features of human and object instances often fails to capture fine-grained contextual cues.
+Inspired by the success of Vision Language Models (VLMs) [Liao2022GENVLKTSA, Lei2024EZHOIVA, Guo2024UnseenNM, Ning2023HOICLIPEK], we leverage CLIP [Radford2021LearningTV] to generate rich multimodal features, thereby enhancing feature interaction between each human-object pair.
+Benefiting from the sufficient prior knowledge of CLIP, our method not only achieves a significant performance gain but also effectively mitigates the performance disparity between rare and non-rare classes, which is a persistent challenge for previous methods [zhang2021scg, zhang2022upt, liao2020ppdm, Ma2023FGAHOIFA].
 Moreover, we introduce a Multi-level Feature Interaction (MFI) mechanism designed to further utilize these multimodal features and enhance the exchange of contextual information.
-This mechanism leverages both low-level and high-level vision and language features to intensify the interaction between human-object pairs, thereby enriching the graph network with more fine-grained semantic information.
-In light of the preceding analysis, we propose Multimodal Graph Network Modeling (MGNM), a novel and efficient two-stage framework for HOI detection.
-Our MGNM achieves state-of-the-art performance on two widely used benchmarks: HICO-DET (Chao et al.
-2018) and V-COCO (Gupta and Malik 2015).
-Furthermore, when integrated with a stronger object detector, our method demonstrates a significant performance gain and maintains an effective balance between rare and non-rare classes.
+Collaborating with the designed object- and interaction-centric prompts, this mechanism leverages both low-level and high-level visual and language features to intensify the feature interaction between human-object pairs, thereby enriching the graph network with more fine-grained semantic information.
+In light of the preceding analysis, we propose Multimodal Graph Network Modeling (MGNM), a novel and effective two-stage framework for HOI detection.
+Our MGNM achieves SOTA performance on two widely used benchmarks: HICO-DET [chao2018learning] and V-COCO [gupta2015visual].
+Furthermore, when integrated with a more advanced object detector, our method demonstrates a significant performance gain and maintains an effective balance between rare and non-rare classes.
 The effectiveness of each proposed component is validated through extensive ablation studies.
@@
   • We propose a novel multimodal graph network framework that explicitly models the relational structure in HOI detection, addressing the limitations of Transformer-based HOI detectors that rely on implicit attention mechanisms.
-  • We design a multi-level feature interaction mechanism that effectively fuses low-level and high-level semantic cues across modalities, substantially enhancing the exchange of fine-grained contextual information within the graph structure.
-  • Extensive experiments on HICO-DET and V-COCO benchmarks demonstrate that our method not only achieves state-of-the-art performance but also maintains an effective balance between rare and non-rare classes.
-Related Work
-One-stage Methods
+  • Collaborating with the object- and interaction-centric prompts, we design a multi-level feature interaction mechanism that effectively fuses low-level and high-level semantic cues across modalities, substantially enhancing the exchange of fine-grained contextual information within the graph structure.
+  • Extensive experiments on HICO-DET and V-COCO benchmarks demonstrate that our method not only achieves SOTA performance but also maintains an effective balance between rare and non-rare classes.
+2 Related Work
+2.1 One-stage Methods
 One-stage methods are characterized by their parallel detection of human and object instances and prediction of their corresponding interactions.
-Early approaches typically relied on heuristics such as human keypoints (Liao et al.
-2020) or predefined union boxes (Kim et al.
-2020).
+Early approaches typically relied on heuristics such as human keypoints [liao2020ppdm, Zhong2021GlanceAG] or predefined union boxes [kim2020uniondet, Gkioxari2017DetectingAR].
 More recently, Transformer-based architectures have become the dominant paradigm for one-stage HOI detection.
 For instance, Chen et al.
-(Chen et al.
-2021) employ two separate Transformer decoders to predict human and object instances and their interactions in parallel.
+[chen_2021_asnet] employ two separate Transformer decoders to predict human and object instances and their interactions in parallel.
 Similarly, Yang et al.
-(Yang et al.
-2023) leveraged VLMs to extract fine-grained semantic features, which were then processed by two sub-decoders to simultaneously predict HOI triplets.
+[yang2023boosting] leveraged VLMs to extract fine-grained semantic features, which were then processed by two sub-decoders to simultaneously predict HOI triplets.
 Nevertheless, one-stage methods are often encumbered by significant memory usage and slow convergence speed.
-Two-stage Methods
+2.2 Two-stage Methods
 Two-stage methods first employ an off-the-shelf object detector to identify human and object instances, and subsequently predict interactions between the matched human-object pairs.
-Recent two-stage methods often use Transformer encoders to extract features or enhance attention mechanisms (Zhang, Campbell, and Gould 2022; Zhang et al.
-2023).
-Another line of work (Guo et al.
-2024; Lei et al.
-2023; Lei, Wang, and Robby T.
-2024) has incorporated VLMs for feature refinement, leading to notable accuracy improvements.
-However, while the Transformer excels as a powerful feature extractor, it does not offer an explicit relational modeling framework for the HOI detection task (Antoun and Asmar 2023).
+Recent two-stage methods often use Transformer to extract features or enhance attention mechanisms [zhang2022upt, zhang2023pvic].
+Another line of work [Guo2024UnseenNM, 11092421, lei2024efficient, NEURIPS2024_2a54def4] has incorporated VLMs for feature refinement, and subsequently adopted Transformer decoders to fuse the features.
+However, while the Transformer excels as a powerful feature extractor, it does not offer an explicit relational modeling framework for the HOI detection task.
 To address this gap, our method employs a multimodal graph network framework instead.
-Our work is situated within a growing line of GNN-based HOI detectors.
-For instance, SCG (Zhang, Campbell, and Gould 2021) utilized a GNN to facilitate information exchange between human and object nodes, while ViPLO (Park, Park, and Lee 2023) incorporated features from the CLIP image encoder and ViTPose (Xu et al.
-2022) to enrich this interaction.
-Nevertheless, a key limitation of these methods is their primary reliance on visual features, which overlooks the rich semantic cues offered by multimodal features, thereby limiting their ultimate performance.
-Method
-Figure 2: Overview of our MGNM framework.The proposed MGNM framework initiates by employing the DETR model to detect human and object instances, which are subsequently matched to form candidate human-object pairs.
-The features of these pairs are then processed through a four-stage multimodal graph network, consisting of the spatial, visual, textual, and interaction stages.
-(1) Spatial Stage: For each candidate pair, low-level spatial features derived from their bounding boxes are used to initialize pairwise representations.
-Simultaneously, these spatial features are utilized to construct a weighted adjacency matrix that regulates feature interactions in the subsequent stages.
+Our work is situated within a growing line of GNN-based HOI detectors [Qi2018LearningHI, Gao2020DRGDR, zhang2021scg, park2023viplo].
+For instance, SCG [zhang2021scg] utilized a GNN to facilitate information exchange between human and object nodes, while ViPLO [park2023viplo] incorporated features from the CLIP image encoder and ViTPose-L [xu2022vitpose] to enrich this feature interaction.
+Nevertheless, a key limitation of these methods is they usually treat GNN as a feature extractor, but don’t analyze it from the perspective of explicit modeling.
+Moreover, they primarily rely on visual features and overlook the rich semantic cues offered by multimodal features, thereby limiting their further performance improvement.
+3 Method
+Figure 2: Overview of our MGNM framework.
+The core of our proposed MGNM framework is a four-stage multimodal graph network.
+(1) Spatial Stage: For each candidate pair, low-level spatial features derived from their bounding boxes and 3D location prior are used to initialize pairwise representations.
+After that, these spatial features are utilized to construct a weighted matrix that regulates feature interactions in the subsequent stages.
 (2) Visual Stage: High-level visual semantic features are extracted using the CLIP image encoder.
-These features further enrich the interactions among human-object pairs by incorporating detailed visual context.
-(3) Textual Stage: The CLIP text encoder is adopted to obtain semantic embeddings for the corresponding human and object categories, providing complementary textual cues for each pair.
-(4) Interaction Stage: In the final stage, the model captures high-level interaction features between human-object pairs, facilitating more effective relational reasoning within the graph structure.
-Finally, the refined pairwise representations are utilized as queries in a Transformer decoder, which attends to global image features from the backbone (serving as keys and values), to predict the final HOI triplets.
+These visual cues further enrich the interactions among human-object pairs.
+(3) Textual Stage: The CLIP text encoder is adopted to obtain semantic cues for the corresponding human and object instances with the designed object-centric prompt.
+(4) Interaction Stage: In the final stage, collaborating with the interaction-centric prompt, the model captures high-level interaction features between human-object pairs, facilitating more effective relational reasoning within the graph structure.
+Finally, the refined pairwise representations are utilized as queries in the Transformer decoder to predict the final HOI triplets.
 This section provides a detailed description of our method.
@@
 An overview of our method is illustrated in Figure 2.
-Following the standard two-stage paradigm for HOI detection, we assume that a given image contains $m$ detected humans and $n$ detected objects, resulting in $m\times n$ candidate human-object pairs.
-Here, we denote $\mathbf{N}\in\mathbb{R}^{(m+n)\times d}$ as the human and object instances, which serves as the node representations in the GNN.
-The corresponding human and object sets within the human-object pairs are represented as $\mathbf{H}\in\mathbb{R}^{(m\times n)\times d}$ and $\mathbf{O}\in\mathbb{R}^{(m\times n)\times d}$, respectively.
+In this paper, we assume a given image contains $n$ detected objects (including $m$ humans).
+With a standard one-to-one matching strategy, we obtain $m\times n$ candidate human-object pairs, which can be seen as the adjacency matrix in GNN.
+Here, we denote $\mathbf{N}\in\mathbb{R}^{n\times d}$ as the human and object instances set, which serves as the node representations in GNN.
+The corresponding human and object sets within the human-object pairs are represented as $\mathbf{H}\in\mathbb{R}^{(m\times n)\times d}$ and $\mathbf{O}\in\mathbb{R}^{(m\times n)\times d}$.
 $d$ is the dimension of the nodes.
 Additionally, the features of human-object pairs are denoted as $\mathbf{E}\in\mathbb{R}^{(m\times n)\times 2d}$, which can be interpreted as the edge embeddings within the GNN framework.
-Multi-level Feature Interaction Mechanism
+3.1 Multi-level Feature Interaction Mechanism
 In this paper, we propose a novel Multi-level Feature Interaction (MFI) mechanism.
-This mechanism utilizes the low-level and high-level vision and language features to enhance the interactive modeling of human-object pairs.
+This mechanism utilizes the low-level and high-level visual and language features to enhance the interactive modeling of human-object pairs.
 As illustrated in Figure 2, the MFI mechanism is structured into four sequential stages: Spatial, Visual, Textual, and Interaction.
@@
 The spatial stage is designed to exploit geometric relationships to enhance interactions between human-object pairs.
-Specifically, the spatial features comprise bounding box coordinates, box areas, Intersection over Union (IoU), and other related metrics (Zhang, Campbell, and Gould 2021).
-These features encode the fundamental geometric configurations of each human-object pair and serve as a source of low-level visual information.
-Moreover, since these are handcrafted features, their computation incurs minimal overhead for the model (Zhang, Campbell, and Gould 2022).
-Upon extracting the spatial features, we utilize a Multimodality Fusion (MMF) mechanism (Zhang et al.
-2023) to integrate them with the initial human and object features, thereby generating the initial representation for each human-object pair.
-$\mathbf{E}$ $=\text{MMF}(\mathbf{H}\oplus\mathbf{O},\mathbf{S}),$ (1)
+Specifically, the spatial features comprise bounding box coordinates, box areas, Intersection over Union (IoU), and other related metrics [zhang2021scg, zhang2022upt].
+More details about spatial features can be found in the supplementary materials.
+These features encode the basic geometric relationships of each human-object pair and serve as a source of low-level visual information.
+Moreover, since these are handcrafted features, their computation incurs minimal overhead for the model [zhang2022upt].
+Upon extracting the spatial features, we utilize a multimodality fusion (MMF) mechanism [zhang2023pvic] to integrate them with the initial human and object features, thereby generating the initial representations for each human-object pair:
+$\mathbf{E}$ $=\text{MMF}\left(\mathbf{H}\oplus\mathbf{O},\mathbf{S}\right),$ (1)
 where $\oplus$ means the concatenation operation and $\mathbf{S}$ means the spatial features.
-In addition, based on the spatial features and corresponding human and object features, the weight matrix $\mathbf{W}_{G}$ for the multimodal graph network is attained with the multi-branch fusion (MBF) mechanism (Zhang, Campbell, and Gould 2021).
-$\mathbf{W}_{G}$ $=\text{Linear}(\text{MBF}(\mathbf{H}\oplus\mathbf{O},\mathbf{S})),$ (2)
-where the linear layer is used to align the dimensions.
+Then, inspired by [11092421], we add the 3D location prior ($\mathbf{L}$) into pair representations.
+As demonstrated by [11092421], these 3D location prior features are beneficial for the model to discriminate between with interaction and no interaction.
+The formulation is as follows:
+$\mathbf{E}$ $=\text{MLP}\left(\mathbf{E}\oplus\mathbf{L}\right).$ (2)
+In addition, based on the spatial features and corresponding human and object features, the weight matrix $\mathbf{W}_{G}\in\mathbb{R}^{m\times n}$ for the multimodal graph network is attained with the multi-branch fusion (MBF) mechanism [zhang2021scg]:
+$\mathbf{W}_{G}$ $=\text{Linear}\left(\text{MBF}\left(\mathbf{H}\oplus\mathbf{O},\mathbf{S}\right)\right),$ (3)
+where the linear layer is used to align the dimensions. More details about the MBF mechanism can be found in the supplementary materials.
 Visual Stage
-The visual stage leverages the CLIP to incorporate high-level vision features.
+The visual stage leverages the CLIP to incorporate high-level visual features.
 Specifically, we employ the CLIP image encoder to extract high-level visual semantic features from the given image.
-These visual features are then used to pass the high-level visual semantic information to the humans and objects as follows:
-$\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}(\mathbf{V},\mathbf{N})$ $=\sigma(\mathbf{W}_{G}\odot\text{MBF}_{v}(\mathbf{V},\mathbf{N})),$ (3)
+These visual features are then used to pass the high-level visual semantic cues to humans and objects as follows:
+$\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}\left(\mathbf{V},\mathbf{N}\right)$ $=\text{ReLU}\left(\mathbf{W}_{G}\odot\text{MBF}_{v}\left(\mathbf{V},\mathbf{N}\right)\right),$ (4)
 where $\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}$ denotes the message passing function responsible for propagating visual information to the human and object node features.
 $\mathbf{V}$ represents the visual features from CLIP.
-$\sigma$ is ReLU activation function and $\odot$ is element-wise multiplication.
-$\text{MBF}_{v}$ is a subclass of the MBF mechanism, tasked with integrating visual and node features.
+$\odot$ is element-wise multiplication.
+$\text{MBF}_{v}$ is a subclass of the MBF mechanism, tasked with integrating visual and node features and providing complementary cues for the nodes.
 Finally, the human and object node features are updated using a layer normalization step, as formulated below:
-$\mathbf{N}$ $=\text{LN}(\mathbf{N}+\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}(\mathbf{V},\mathbf{N})),$ (4)
+$\mathbf{N}$ $=\text{LN}\left(\mathbf{N}+\text{M}_{\mathcal{V}\rightarrow\mathcal{N}}\left(\mathbf{V},\mathbf{N}\right)\right),$ (5)
 where LN means the layer normalization.
@@
 The textual stage is dedicated to generating textual embeddings for each node, providing a source of low-level language features.
-As illustrated in Figure 2, we adopt the prompt template “a photo of a $\langle\text{object}\rangle$”, where the $\langle\text{object}\rangle$ comes from the detection results.
+As illustrated in Figure 2, we adopt a simple object-centric prompt template “a $\langle\text{object}\rangle$ in the human-centered scene”, focusing on the expected object.
+The $\langle\text{object}\rangle$ comes from the detection results.
 In this paper, for brevity, we denote the humans and objects as the whole node set $\mathbf{N}$, and the corresponding textual features are assumed to be matched.
 The process is formulated as follows:
-$\text{M}_{\mathcal{T}\rightarrow\mathcal{N}}(\mathbf{T},\mathbf{N})$ $=\sigma(\mathbf{W}_{G}\odot\text{MBF}_{t}(\mathbf{T},\mathbf{N})),$ (5)
+$\text{M}_{\mathcal{T}\rightarrow\mathcal{N}}(\mathbf{T},\mathbf{N})$ $=\text{ReLU}\left(\mathbf{W}_{G}\odot\text{MBF}_{t}\left(\mathbf{T},\mathbf{N}\right)\right),$ (6)
 where $\text{M}_{\mathcal{T}\rightarrow\mathcal{N}}$ means the message passing function responsible for propagating textual information to the human and object node features.
@@
 $\text{MBF}_{t}$ is a subclass of the MBF mechanism.
-Similarly, we also use layer normalization step when updating the human and object features.
+Similarly, we also use the layer normalization step when updating the human and object features.
 $\mathbf{N}$ $=\text{LN}(\mathbf{N}+\text{M}_{\mathcal{T}\rightarrow\mathcal{N}}(\mathbf{T},\mathbf{N})).$ (7)
 Interaction Stage
-To incorporate high-level interaction semantic cues, we again employ the CLIP text encoder.
+To incorporate high-level interaction semantic cues, we propose the interaction-centric prompt.
+Specifically, this prompt is constructed using geometric and spatial cues closely related to the interaction, including IoU, relative size, relative position, and center distance.
+Based on the specific values of these cues, we dynamically generate distinct prompts for each human-object pair.
+For IoU, we define four descriptive phrases: “closely interacting with”, “directly engaging with”, “potentially interacting with”, and “near but may not be interacting with”.
+Similarly, for relative size, descriptive terms such as “very large”, “large”, and “small” are employed.
+By combining these dynamically selected phrases, the complete interaction-centric prompt is formed.
+For instance, an $\text{IoU}>0.3$ triggers the phrase “closely interacting with”, while a $\text{relative size}>3.0$ appends “very large $\langle\text{object}\rangle$”, where $\langle\text{object}\rangle$ is the predicted object class.
+The resulting prompt becomes: “a person closely interacting with very large $\langle\text{object}\rangle$”.
+In this situation, given the high IoU implying a direct interaction, explicit position and distance cues are deemed unnecessary.
+Conversely, $\text{IoU}\leq 0.05$, yields the phrase “near but may not be interacting with”.
+If this is combined with a $\text{relative size}\leq 0.3$, the prompt “very small $\langle\text{object}\rangle$” is added.
+Subsequently, a $\text{center distance}\in(0.15,0.3)$ appends positional descriptors such as “right” and “above”.
+Finally, the complete prompt becomes: “a person near but may not be interacting with very small $\langle\text{object}\rangle$ positioned above and to the right”.
+A comprehensive definition of these prompt-generation rules can be found in the supplementary materials.
+This interaction-centric prompt design provides discriminative semantic cues for the text encoder.
+Leveraging this interaction-centric prompt, we once again employ the CLIP text encoder.
 This time, it is used to generate interaction-centric semantic features, which constitute high-level language features.
-The prompt template “a photo of a person interacting with $\langle\text{object}\rangle$” is adopted, which aims to boost the feature interaction between human-object pairs.
 This process is formulated as follows:
-$\mathbf{E}$ $=\text{LN}(\mathbf{E}+\mathbf{I}),$ (7)
+$\mathbf{E}$ $=\text{LN}\left(\mathbf{E}+\mathbf{I}\right),$ (8)
 where $\mathbf{I}$ denotes the interaction features of the human-object pairs.
 As a preliminary step, the human-object pair features are first enriched by integrating the features of their corresponding human and object nodes as follows:
-$\mathbf{E}$ $=\text{LN}(\mathbf{E}+\mathbf{H}\oplus\mathbf{O}).$ (8)
-Multimodal Graph Network Modeling
+$\mathbf{E}$ $=\text{LN}\left(\mathbf{E}+\mathbf{H}\oplus\mathbf{O}\right).$ (9)
+3.2 Multimodal Graph Network Modeling
 Building upon the MFI mechanism detailed above, this section presents the complete architecture of the proposed MGNM framework.
@@
 At the core of our framework lies the MFI mechanism, as detailed in Algorithm 1 (Lines 1–7).
-The process commences by constructing the initial human-object pair features, which are then iteratively refined through the proposed four-stage pipeline.
-Following (Park, Park, and Lee 2023), we set the number of iterative steps $T$ to 2.
+The process commences by constructing the initial human-object pair representations, which are then iteratively refined through the proposed four-stage pipeline.
+Following [park2023viplo], we set the number of iterative steps $T$ to 2.
 Notably, in the visual and textual stages, we leverage visual and textual features to update the human and object instances, thereby infusing multimodal information into the node encodings.
-In addition, since the message passing is guided by a weight matrix computed from the relational structure among human-object pairs, this implicitly promotes feature interaction between pairs.
+In addition, since the message passing is guided by a weight matrix computed from the relational structure among human-object pairs, these two stages implicitly promote feature interaction between pairs.
 In contrast, the spatial and interaction stages are explicitly designed to enhance feature communication between human-object pairs.
@@
 In this decoder, the human-object pair features act as queries, while the global features extracted from the backbone serve as keys and values.
-Experiments
+4 Experiments
 HICO-DET V-COCO
 Method Backbone Default Known Object
-Full Rare Non-Rare Full Rare Non-Rare $AP^{S1}_{role}$ $AP^{S2}_{role}$
+Full Rare Non-Rare Full Rare Non-Rare $\text{AP}^{S1}_{role}$ $\text{AP}^{S2}_{role}$
 One-stage methods
-UnionDet (2020) R50-F 17.58 11.72 19.33 19.76 14.68 21.27 47.5 56.2
-PPDM (2024) Swin-B-F 27.59 18.07 30.44 29.11 18.85 32.71 - -
-AS-Net (2021) R50 28.87 24.25 30.25 31.74 27.07 33.14 53.9 -
-FGAHOI-S (2024b) Swin-T 29.94 22.24 32.24 32.48 24.16 34.97 60.5 61.2
-DOQ (2022) R50+C 33.28 29.19 34.50 - - - 63.5 -
-GEN-VLTK (2022) R50+C 33.75 29.25 35.10 36.78 32.75 37.99 62.4 64.4
-DiffHOI-S (2023) R50+C+SD 34.41 31.07 35.40 37.31 34.56 38.14 61.1 63.5
-FGAHOI-L (2024b) Swin-L 37.18 30.71 39.11 38.93 31.93 41.02 - -
-DiffHOI-L (2023) Swin-L+C+SD 40.63 38.10 41.38 43.14 40.24 44.01 65.7 68.2
+UnionDet [kim2020uniondet] R50 17.58 11.72 19.33 19.76 14.68 21.27 47.5 56.2
+AS-Net [chen_2021_asnet] R50 28.87 24.25 30.25 31.74 27.07 33.14 53.9 -
+FGAHOI [10315071] Swin-T 29.94 22.24 32.24 32.48 24.16 34.97 60.5 61.2
+DOQ [qu2022distillation] R50+CLIP 33.28 29.19 34.50 - - - 63.5 -
+DiffHOI [yang2023boosting] R50+CLIP+SD 34.41 31.07 35.40 37.31 34.56 38.14 61.1 63.5
+MP-HOI [Yang2024OpenWorldHI] R50+CLIP+SD 36.50 35.48 36.80 - - - 66.2 67.6
+FGAHOI-L [10315071] Swin-L 37.18 30.71 39.11 38.93 31.93 41.02 - -
+DiffHOI-L [yang2023boosting] Swin-L+CLIP+SD 40.63 38.10 41.38 43.14 40.24 44.01 65.7 68.2
+$\text{MP-HOI-L}^{\dagger}$ [Yang2024OpenWorldHI] Swin-L+CLIP+SD 44.53 44.48 44.55 - - - - -
 Two-stage methods
-SCG (2021) R50-F 29.26 24.61 30.65 32.87 27.89 34.35 54.2 60.9
-PPDM++ (2024) Swin-B-F 30.10 23.73 32.00 31.80 24.93 33.85 - -
-UPT (2022) R101 32.31 28.55 33.44 35.65 31.60 36.86 60.7 66.2
-ADA-CM (2023) R50+C 33.80 31.72 34.42 - - - 60.7 66.2
-ILCN (2025) Swin-T 33.80 29.83 34.99 - - - 63.4 65.3
-PViC w/ DETR (2023) R50 34.69 32.14 35.45 38.14 35.38 38.97 62.8 67.8
-HOIGen (2024) R50+C 34.84 34.52 34.94 - - - - -
-ViPLO (2023) R101-D+C+VP 34.95 33.83 35.28 38.15 36.77 38.56 60.9 66.6
-TKCE (2025) R50+C 35.11 34.35 35.34 - - - - -
-LAIN (2025) R50+C 36.02 35.70 36.11 - - - 63.4 65.3
-MP-HOI-S (2024) R50+C+SD 36.50 35.48 36.80 - - - 66.2 67.6
-PViC w/ $\mathcal{H}$-DETR (2023) Swin-L 44.32 44.61 44.24 47.81 48.38 47.64 64.1 70.2
-MP-HOI-L (2024) Swin-L+C+SD 44.53 44.48 44.55 - - - - -
-MGNM R50+C 36.68 34.10 37.45 39.99 37.15 40.84 62.1 67.5
-MGNM w/ $\mathcal{H}$-DETR Swin-L+C 46.54 48.09 46.07 49.78 50.82 49.47 64.3 70.5
+SCG [zhang2021scg] R50 29.26 24.61 30.65 32.87 27.89 34.35 54.2 60.9
+PPDM++ [10496247] Swin-B 30.10 23.73 32.00 31.80 24.93 33.85 - -
+UPT [zhang2022upt] R101 32.31 28.55 33.44 35.65 31.60 36.86 60.7 66.2
+ILCN [lu2025intra] Swin-T 33.80 29.83 34.99 - - - 63.4 65.3
+PViC [zhang2023pvic] R50 34.69 32.14 35.45 38.14 35.38 38.97 62.8 67.8
+HOIGen [Guo2024UnseenNM] R50+CLIP 34.84 34.52 34.94 - - - - -
+ViPLO [park2023viplo] R101-DC+CLIP+ViTPose 34.95 33.83 35.28 38.15 36.77 38.56 60.9 66.6
+TKCE [10889833] R50+CLIP 35.11 34.35 35.34 - - - - -
+LAIN [kim2025locality] R50+CLIP 36.02 35.70 36.11 - - - 63.4 65.3
+VRDiff [Cao_2025_ICCV] R50+CLIP+SD 36.77 35.66 37.11 39.98 38.69 40.36 - -
+$\text{ADA-CM}^{\dagger}$ [ting2023hoi] R50+CLIP 38.40 37.52 38.66 - - - 58.6 64.0
+RLIPv2 [Yuan2023RLIPv2FS] Swin-T+BLIP 38.60 33.66 40.07 - - - 68.8 70.8
+HORP [11092421] R50+CLIP+Gaze 38.61 36.14 39.34 40.98 38.25 41.79 65.6 68.3
+$\text{HOLa}^{\dagger}$ [lei2025hola] R50+CLIP+GPT 39.05 38.66 39.17 - - - - -
+PViC-L [zhang2023pvic] Swin-L 44.32 44.61 44.24 47.81 48.38 47.64 64.1 70.2
+Pose-aware [wu2024exploring] Swin-L+ViTPose 46.01 46.74 45.80 49.50 50.59 49.18 63.0 68.7
+HORP-L [11092421] Swin-L+CLIP+Gaze 47.53 46.81 47.74 51.24 50.78 51.38 68.9 71.1
+MGNM R50+CLIP 39.25 38.66 39.42 42.36 41.76 42.54 67.7 70.2
+$\text{MGNM}_{L}^{\dagger}$ Swin-L+CLIP 49.09 50.62 48.63 52.78 53.89 52.45 70.4 73.9
 Table 1: Performance comparison in terms of mAP (%) on the HICO-DET and V-COCO dataset.
-To ensure a fair comparison, we report both the performance and the corresponding feature backbone for each method.
 In each evaluation metric, the best result is marked in bold and the second-best result is underlined.
-In this paper, R50, R101, R50-F and R101-D denotes ResNet50, ResNet101, ResNe50-FPN and ResNet101-DC5 model, respectively.
-A similar notation applies to Swin Transformer (Liu et al.
-2021).
-C, SD, and VP mean CLIP, Stable Diffusion, and ViTPose models, respectively.
-We note that MP-HOI (Yang et al.
-2024) utilizes additional training data (Magic-HOI and SynHOI), which should be considered when comparing results.
+In this paper, R50, R101, and R101-DC denotes ResNet50, ResNet101, and ResNet101 with dilated convolution, respectively.
+SD means Stable Diffusion model [Rombach2021HighResolutionIS].
+$\dagger$ means the CLIP adopts the ViT-L image encoder [Dosovitskiy2020AnII].
+Others adopt the ViT-B image encoder.
 This section presents a series of experiments designed to evaluate the effectiveness of our proposed method.
 First, we briefly introduce the experimental settings.
-Next, we present a comprehensive comparison of our method against existing state-of-the-art methods.
-Subsequently, the ablation study of our method is conducted.
-Finally, we rethink and analyze the Rare-Non-rare bias in the HOI detection field and provide some promising suggestions.
-Experimental Settings
-Datasets
+Next, we present a comprehensive comparison of our method against the SOTA methods.
+Subsequently, we conduct the ablation study of our method.
+All of its experiments are conducted on the HICO-DET dataset under Deafult setting.
+Finally, we analyze the Rare-Non-rare bias in the HOI detection field and provide some promising suggestions.
+4.1 Experimental Settings
+Datasets and Evaluation Metrics
 Our experiments are conducted on two widely used public benchmarks: HICO-DET and V-COCO.
-HICO-DET contains 38,118 training images and 9,658 test images.
-It includes 80 object classes from the MS COCO dataset (Lin et al.
-2014) and 117 action classes, which combine to form 600 types of HOI instances.
-V-COCO, a smaller-scale subset of the MS COCO dataset, contains 2,533 training, 2,867 validation, and 4,946 test images, encompassing 80 object classes and 29 action classes.
+Following the standard evaluation protocols in [zhang2023pvic, park2023viplo], we report the mean average precision (mAP) for evaluation.
 Given its larger scale and closer alignment with real-world data distributions, our primary experimental focus is on the HICO-DET dataset.
-Evaluation Metrics
-Following the standard evaluation protocols in (Park, Park, and Lee 2023), we report the mean Average Precision (mAP) for evaluation.
-An HOI triplet is considered a true positive when the IoU of the detected human and object is larger than 0.5 and the action is predicted accurately.
-For the HICO-DET dataset, we report performance under two standard evaluation settings: Default and Known Object.
-Each setting has three different subsets: (1) all 600 HOI classes (Full), (2) 138 HOI classes with less than 10 training instances (Rare), and (3) 462 HOI classes with 10 or more training instances (Non-Rare).
-For the V-COCO dataset, we evaluate performance based on two scenarios: Scenario 1, which requires correct prediction of occluded object bounding boxes, and Scenario 2, which does not.
+More details about datasets and evaluation metrics can be found in the supplementary materials.
 Implementation Details
-For HICO-DET and V-COCO datasets, we use the fine-tuned DETR and $\mathcal{H}$-DETR (Jia et al.
-2023) weights provided by Zhang et al.
-(Zhang et al.
-2023).
+For HICO-DET and V-COCO datasets, following [zhang2023pvic, 11092421], we use the fine-tuned DETR and $\mathcal{H}$-DETR [Jia2022DETRsWH] models as the object detector.
 During training, following the commonly used training paradigm of two-stage methods, the weights of the object detector are frozen, while only the interaction predictor is trained.
-Meanwhile, the focal loss (Lin et al.
-2017) is employed as the loss function.
-Following (Lei et al.
-2023), we also incorporate the adapter into the CLIP module.
+Meanwhile, the focal loss [Lin2017FocalLF] is employed as the loss function.
+Following [ting2023hoi, lei2025hola], we employ the CLIP ViT-B/16 and ViT-L/14@336px models for our method, and all the parameters of them are frozen, except for the adapter layer.
+We also apply data augmentation techniques from previous works [ting2023hoi, zhang2023pvic], including random cropping, resizing, and color jittering.
+During inference, following [zhang2023pvic, zhang2022upt], we use the hyparameter $\lambda$ to suppress overconfident objects.
 The model is optimized using the AdamW optimizer with an initial learning rate of $10^{-4}$.
-All models are trained for 30 epochs on 4 NVIDIA A100 GPUs, using a total batch size of 32.
-Unless otherwise specified, all other hyperparameters adhere to the settings established in (Zhang et al.
-2023; Lei et al.
-2023).
-Comparison with State-of-the-art Methods
-Table 1 provides a comprehensive performance comparison between our proposed MGNM and current state-of-the-art methods.
+All models are trained for 30 epochs on 4 NVIDIA A100 GPUs, using a batch size of 16.
+4.2 Comparison with State-of-the-art Methods
+Table 1 provides a comprehensive performance comparison between our proposed MGNM and current SOTA methods.
 For clarity, the competing methods are grouped into one-stage and two-stage categories.
 On the HICO-DET dataset, MGNM outperforms all other one-stage and two-stage approaches that utilize a ResNet backbone.
-Specifically, in the Default Full setting, our method achieves a 2.27 mAP improvement over DiffHOI-S, despite the latter leveraging two VLMs, namely CLIP and Stable Diffusion (Rombach et al.
-2021).
-MGNM also surpasses the leading two-stage approach, MP-HOI-S, by 0.18 mAP in the Default Full setting, even though MP-HOI-S incorporates both additional VLMs and extra training data.
-Moreover, MGNM exceeds the performance of the recent GNN-based ViPLO by 1.73 mAP in the Default Full setting, notwithstanding ViPLO’s adoption of the supplementary ViTPose-L model (Xu et al.
-2022).
+Specifically, in the Default Full setting, our method achieves a 4.84 and 2.75 mAP improvement over the Transformer-based methods DiffHOI and MP-HOI, despite they leveraging two VLMs, namely CLIP and Stable Diffusion.
+MGNM also surpasses the leading two-stage approaches, HORP and HOLa, by 0.64 and 0.2 mAP in the Default Full setting, even though they incorporates additional models.
+Moreover, MGNM exceeds the performance of the recent GNN-based method, ViPLO, by 4.3 mAP in the Default Full setting, notwithstanding ViPLO’s adoption of the supplementary ViTPose-L model [xu2022vitpose].
 In the Known Object setting, MGNM consistently outperforms all ResNet-based counterparts across all three subsets by a substantial margin.
-Notably, our ResNet50-based MGNM even surpasses FGAHOI-L, which utilizes a more powerful Swin-L backbone, by 1.06 mAP in the Full setting.
+Notably, our R50-based MGNM even surpasses FGAHOI-L, which utilizes a more powerful Swin-L backbone [Liu2021SwinTH], by 3.43 mAP in the Full setting.
 Collectively, these results demonstrate the effectiveness of our multimodal graph network and its explicit modeling strategy for HOI detection.
-On the V-COCO dataset, MGNM also achieves highly competitive performance among ResNet-based methods.
-We note that MP-HOI-S obtains a higher $AP_{role}^{S1}$, which we attribute to its use of additional training data.
+On the V-COCO dataset, MGNM also achieves substantial performance improvement among ResNet-based methods.
+We note that RLIPv2 obtains a higher performance on V-COCO, which we attribute to its large-scale pre-training on additional data.
 This observation suggests that GNN-based approaches, including MGNM, could further benefit from large-scale pre-training, highlighting a promising direction for future research.
+To further demonstrate the scalability of our approach, we compare the performance of MGNM equipped with a more advanced object detector against other SOTA methods that utilize similarly strong backbones.
+As shown in Table 1, $\text{MGNM}_{L}$ achieves substantial performance improvements over the competing methods.
+For example, on the HICO-DET Default Full setting, our method outperforms the Transformer-based methods, PViC-L and Pose-aware, by 4.77 and 3.08 mAP.
+Notably, although both DiffHOI-L and MP-HOI-L leverage two VLMs each, our single-VLM approach achieves significant gains of 8.46 and 4.56 mAP over these methods, respectively.
+These results underscore not only the scalability of our framework but also the effectiveness of our multimodal graph network design.
 Methods Full Rare Non-Rare
-w/o Spatial Stage 36.00 34.68 36.39
-w/o Visual Stage 35.32 32.66 36.12
-w/o Textual Stage 36.16 33.81 36.86
-w/o Interaction Stage 35.61 33.37 36.28
-vanilla 36.68 34.10 37.45
-Table 2: Effectiveness of each component within our MFI mechanism. w/o means “without”.
-Furthermore, to further demonstrate the scalability of our approach, we compare the performance of MGNM equipped with a more powerful object detector against other state-of-the-art methods that utilize similarly strong backbones.
-As shown in Table 1, MGNM integrated with $\mathcal{H}$-DETR achieves substantial performance improvements over the competing methods.
-For example, on the HICO-DET Default Full setting, our method outperforms PViC with $\mathcal{H}$-DETR by 2.22 mAP.
-Notably, although both DiffHOI-L and MP-HOI-L leverage two VLMs each, our single-VLM approach achieves significant gains of 5.91 and 2.01 mAP over these methods, respectively.
-These results underscore not only the scalability of our framework but also the effectiveness of our multimodal graph network design.
-Ablation Study
+w/o Spatial Stage 38.17 37.62 38.33
+w/o Visual Stage 36.78 35.64 37.12
+w/o Textual Stage 38.23 37.44 38.47
+w/o Interaction Stage 36.94 35.18 37.47
+Simple Prompt 37.53 35.92 38.01
+MLP Fusion 37.08 35.16 37.65
+vanilla 39.25 38.66 39.42
+Table 2: Effectiveness of each component within our framework. w/o means “without”.
+4.3 Ablation Study
 (a) carry sports ball
-(b) ride snowboard
-(c) drive boat
-(d) Visualization for c
-(e) Visualization for c
-Figure 3: Qualitative results on the HICO-DET test set.
-Figures 3(a) and 3(b) illustrate successful predictions across a variety of challenging scenarios.
-In contrast, Figure 3(c) presents a representative failure case, wherein the model generates a false positive triplet.
-Specifically, it incorrectly infers that both individuals are driving the boat, whereas only the person on the right is performing this action.
-To further investigate this failure case, Figures 3(d) and 3(e) visualize the corresponding attention maps, highlighting the model’s regions of focus that led to this prediction.
-Methods Backbone mAP $\uparrow$ Time $\downarrow$ Memory $\downarrow$
-SCG R101 29.26 182 ms 7,702 MiB
-ViPLO R101-D+C+VP 34.95 166 ms 10,224 MiB
-MGNM R50+C 36.68 125 ms 2,946 MiB
-Table 3: Comparison of mAP, inference time, and GPU memory on the HICO-DET test set.
-We select two prominent GNN-based baselines for this comparison: SCG and ViPLO.
-It is important to note that the official implementations of both SCG and ViPLO execute object detection and interaction prediction as two separate, sequential processes.
-Therefore, to ensure a fair comparison, our reported time and memory for these baselines encompass the complete process.
-All experiments are conducted with a batch size of 1 on a single NVIDIA A100 GPU.
-Time refers to the elapsed time required for processing one image.
-In this subsection, we conduct ablation studies to analyse the effectiveness of our proposed components and the efficiency of our MGNM framework.
-All the experiments are conducted on the HICO-DET test set under Deafult setting.
+(b) carry backpack
+(c) sit on bench
+(d) wear tie
+(e) ride snowboard
+(f) drive boat
+(g) Visualization for f
+(h) Visualization for f
+Figure 3: Qualitative results on the HICO-DET dataset.
+Figures 3(a)-3(e) illustrate successful predictions across a variety of challenging scenarios.
+Figure 3(f) presents a representative failure case, and Figures 3(g) and 3(h) visualize its corresponding attention maps.
+Methods Full Rare Non-Rare
+CLIP-ViT-B/32 38.83 38.24 39.01
+CLIP-ViT-B/16 39.25 38.66 39.42
+CLIP-ViT-L/14 42.97 42.32 43.17
+CLIP-ViT-L/14@336px 43.53 42.88 43.73
+Table 3: Comparison of our method with different CLIP sizes.
 Components Analysis
 We first analyze the contribution of each of the four stages within our MFI mechanism, which collectively process low-level and high-level multimodal features.
-As shown in Table 2, removing the low-level feature stages, namely the Spatial (row 1) and Textual (row 3) stages, results in performance drops of 0.68 and 0.52 mAP, respectively.
-This demonstrates that these low-level spatial and textual features provide a valuable supplement to the foundational human-object pair representations.
-Ablating the high-level feature stages, the Visual (row 2) and Interaction (row 4) stages, induces a more substantial performance degradation, with drops of 1.36 and 1.07 mAP, respectively.
+As shown in Table 2, in row 1 (w/o Spatial Stage), we only ablate the spatial features, preserving the 3D location prior and weight matrix.
+This results in performance drops of 1.08 mAP.
+In row 3 (w/o Textual Stage), it leads to a performance drop of 1.02 mAP.
+This phenomenon demonstrates that these low-level spatial and textual features provide a valuable supplement to the foundational human-object pair representations.
+Ablating the high-level feature stages, the Visual (row 2) and Interaction (row 4) stages, induces a more substantial performance degradation, with drops of 2.47 and 2.31 mAP, respectively.
 These decreases are considerably larger than those observed in low-level stages.
-This finding suggests that the high-level features are more critical to our model’s predictive power, an observation that aligns with the notion that GNNs are inherently biased towards modeling high-level structural information (Antoun and Asmar 2023).
-Interestingly, we observe a counter-intuitive phenomenon in row 1: ablating the spatial stage improves performance on the rare classes.
-This suggests that low-level vision features may inadvertently amplify the model’s existing bias towards more frequent, non-rare classes.
-Nevertheless, considering their substantial positive impact on the Full and Non-Rare settings, these spatial features remain an indispensable component for overall model performance.
-Efficiency Analysis
-This subsection presents an efficiency analysis of MGNM in comparison to two related GNN-based methods, focusing on inference time and GPU memory usage.
-As reported in Table 3, both SCG and ViPLO decouple the object detection and interaction prediction processes, whereas our method integrates these stages into a unified end-to-end pipeline.
-To ensure a fair comparison, we report the total inference time and GPU memory for SCG and ViPLO, excluding file I/O operations.
-As shown in Table 3, our approach not only achieves substantially higher mAP but also demonstrates superior efficiency, recording the lowest inference time and GPU memory consumption among the compared methods.
-These results highlight the strong balance between effectiveness and efficiency achieved by MGNM.
-It is worth noting that the official object detection scripts provided by the baseline methods appear to be unoptimized with respect to memory usage.
-When replacing their detection modules with DETR weights, resulting in less GPU consumption than the actual GPU memory usage, the measured occupancy drops to 1,904 MiB for SCG and 4,426 MiB for ViPLO.
-Even under these conditions, when comprehensively considering both mAP and inference time, our method continues to deliver the best overall effectiveness and efficiency.
+This finding suggests that the high-level features are more critical to our model’s predictive power, an observation that aligns with the notion that GNNs are inherently biased towards modeling high-level structural information [antoun2023human].
+Next, we analyze the other two important components in our method.
+In row 5 (Simple Prompt), to analyze the effectiveness of the interaction-centric prompt, we replace the interaction-centric prompt with a simple prompt “a photo of person interacting with $\langle\text{object}\rangle$”.
+As shown in Table 2, it results in a performance drop of 1.72 mAP.
+This demonstrates the effectiveness of our interaction-centric prompt, which provides more fine-grained semantic cues for the text encoder.
+At last, in row 6 (MLP Fusion), we ablate the GNN structure in our method and only use a simple MLP fusion to fuse the corresponding multimodal features for the instances.
+This can be interpreted as a non-GNN baseline.
+It can be observed that this ablation results in a drop of 2.17 mAP.
+This demonstrates the significance of our explicit modeling strategy.
+Based on the above analysis, the systematic ablation studies collectively demonstrates the effectiveness of our explicit multimodal graph network modeling.
+Analysis on CLIP Size
+CLIP, as a classic vision language model, has been widely adopted in HOI detection and demonstrated an effective model [ting2023hoi, Lei2024EZHOIVA, park2023viplo].
+Since different versions of CLIP differ significantly in model architecture and parameter count, in this subsection, we analyze the influence of different CLIP model scales on our method.
+As shown in Table 3, we find that CLIP with the same architecture achieves competitive performance.
+For instance, CLIP-ViT-B/16 outperforms CLIP-ViT-B/32 by 0.42 mAP in the Full setting.
+This small margin is expected.
+While ViT-B/16 produces more detailed visual features, they share the same base architecture, and the parameter count is almost same.
+Furthermore, other modules in our HOI detector also provide fine-grained features, which likely further mitigates the gap between the ViT-B/32 and ViT-B/16 variants.
+In contrast, a significant difference emerges between the ViT-B and ViT-L model families.
+For example, CLIP-ViT-L/14 achieves a significant improvement of 3.72 mAP over CLIP-ViT-B/16.
+This is expected, as ViT-L/14 has a larger architecture and possesses approximately $3.5\times$ as many parameters as ViT-B/16, resulting in a much larger model capacity.
+Additionally, the text encoders in different CLIP sizes are architecturally similar, varying primarily in their projection dimension.
+Consequently, the improvement is primarily attributable to the more powerful visual encoder.
+Based on this analysis, adopting a larger CLIP model appears to be an effective strategy for improving performance in HOI detection.
+We also conduct ablation studies on CLIP sizes with the $\mathcal{H}$-DETR.
+Although the performance gain is not as significant as that with R50, it still brings an effective improvement of 1.17 mAP.
+More discussion about this can be found in the supplementary materials.
 Figure 4: Comparison of our method and two related GNN-based methods on rare and non-rare classes.
-Rethinking Rare-Non-rare Bias
+4.4 Rare-Non-rare Bias
 The challenge of few-shot learning is particularly pronounced in HOI detection, where it manifests as a significant performance disparity between rare and non-rare classes.
 As evidenced in Table 1, this long-tail distribution problem is severe in many methods.
-For instance, PPDM and FGAHOI-S exhibit a performance gap of over 10 mAP between their rare and non-rare classes.
+For instance, PPDM and FGAHOI exhibit a performance gap of over 10 mAP between their rare and non-rare classes.
 Similarly, in Figure 4, the GNN-based method SCG also has a significant difference between rare and non-rare classes.
@@
 This suggests that the rich prior knowledge in VLMs can effectively alleviate the Rare-Non-rare bias.
-Our own results, illustrated in Figure 4, support this hypothesis: both our MGNM and ViPLO, which leverage CLIP, demonstrate a much narrower performance gap between rare and non-rare classes.
-Interestingly, ViPLO achieves a better balance than our method.
-We hypothesize that its use of an additional ViTPose-L model provides another source of rich prior knowledge, playing a similar role like VLMs.
-Moreover, as shown in Table 1 and Figure 4, an unexpected trend emerges for methods equipped with the $\mathcal{H}$-DETR object detector, these models exhibit a remarkable reversal, where performance on rare classes actually exceeds that on non-rare classes.
-We posit that this is likely attributable to the unique architectural design of $\mathcal{H}$-DETR, and a deeper investigation into this phenomenon presents a compelling direction for future research.
-In summary, our analysis indicates that leveraging VLMs and advanced object detectors represents a highly effective strategy for mitigating the Rare-Non-rare bias in HOI detection.
-Qualitative Results and Analysis
+Results illustrated in Figure 4 also support this hypothesis: both MGNM and ViPLO, which leverage CLIP, demonstrate a much narrower performance gap between rare and non-rare classes.
+Moreover, as shown in Table 1 and Figure 4, an unexpected trend emerges for methods equipped with a more advanced object detector, such as MP-HOI-L, PViC-L, Pose-aware and our $\text{MGNM}_{L}$.
+These models exhibit a remarkable reversal, where performance on rare classes actually exceeds that on non-rare classes.
+We posit that this is attributable to the unique architectural design of DDETR [Zhang2022DINODW] or $\mathcal{H}$-DETR with Swin-L backbone, and a deeper investigation into this phenomenon presents a compelling direction for future research.
+In summary, our analysis indicates that leveraging VLMs and advanced object detectors represents a highly effective strategy for mitigating the Rare-Non-rare bias.
+5 Qualitative Results and Limitations
 This section presents a qualitative analysis of our method’s predictions, highlighting both its strengths and current limitations.
-As illustrated in Figures 3(a) and 3(b), our model demonstrates robustness by accurately predicting HOI triplets even when the target object is small or occluded.
-Conversely, Figure 3(c) presents a representative failure case involving a false positive prediction.
+As illustrated in Figures 3(a)-3(e), our model demonstrates robustness by accurately predicting HOI triplets in single-person or multi-person scenarios, even when the target object is small or occluded.
+Conversely, Figure 3(f) presents a representative failure case involving a false positive prediction.
 Here, the model incorrectly infers that two individuals are driving the boat, a prediction that contradicts common sense in the real world.
-To diagnose this error, we visualize the corresponding attention maps in Figures 3(d) and 3(e).
+To diagnose this error, we visualize the corresponding attention maps in Figures 3(f) and 3(h).
 The visualization reveals that the model correctly identifies promising regions associated with the drive action for both individuals.
-However, it lacks the commonsense knowledge to understand that only one person can drive a single boat, leading to the erroneous prediction.
+However, it lacks common-sense knowledge to understand that only one person can drive a single boat, leading to the erroneous prediction.
 Furthermore, attention is diffuse and does not pinpoint fine-grained cues, such as hand-on-wheel placement.
-This limitation highlights a crucial direction for future work: developing methods to inject real-world commonsense reasoning into HOI detectors to resolve such logical ambiguities.
-Conclusion
-In this paper, we introduce Multimodal Graph Network Modeling (MGNM), a novel and effective two-stage HOI detection framework.
-The central contribution of MGNM lies in its explicit modeling of the HOI detection task via the construction of a multimodal graph network.
-Furthermore, by leveraging Vision Language Models (VLMs), we design a multi-level feature interaction mechanism that systematically fuses low-level and high-level vision and language features, thereby enhancing information propagation within the graph structure.
-Extensive experiments validate that our approach not only achieves state-of-the-art performance but also delivers superior computational efficiency.
-In addition, our analysis of the pervasive class imbalance issue provides new insights and practical recommendations for mitigating the Rare-Non-rare bias in future research.
+This limitation highlights a crucial direction for future work: developing methods with real-world common-sense reasoning to resolve such logical ambiguities.
+6 Conclusion
+In this paper, we introduce multimodal graph network modeling, a novel and effective two-stage HOI detection framework.
+The central contribution of MGNM lies in its explicit modeling of the HOI detection task via a multimodal graph network.
+Furthermore, by leveraging vision language models and two designed prompts, we design a multi-level feature interaction mechanism that systematically fuses low-level and high-level visual and language features, thereby enhancing information propagation within the graph structure.
+Extensive experiments validate that our approach achieves the SOTA performance.
+In addition, our analysis of the pervasive class imbalance issue provides new insights and practical suggestions for mitigating the Rare-Non-rare bias in future research.